OpenStudy (lina17):

4/5 x + 4/3 = 2x help step by step please

OpenStudy (lina17):

\[\frac{ 4 }{ 5} x+\frac{ 4 }{ 3 }=2x\]

OpenStudy (anonymous):

multiply everything by 15 to get 12x+20x=30x

OpenStudy (anonymous):

i mean 12x+20=30x oops

OpenStudy (anonymous):

subtract 12x to get 20=18x

OpenStudy (anonymous):

i got x=10/9

OpenStudy (anonymous):

that is correct

OpenStudy (lina17):

how did you get that

OpenStudy (anonymous):

idk i might be wrong

OpenStudy (anonymous):

20=18x divide by 18 to get: 20/18=x or 10/9=x

OpenStudy (lina17):

yea but how did you get 20=18x

OpenStudy (anonymous):

look at my previous work to see how I got to 20=18x

jhonyy9 (jhonyy9):

@Bill_Cipher and @sarina.lara just step by step - @lina17 not have asked for answer directly he wan learning how need solving same exercises - ok ?

jhonyy9 (jhonyy9):

@lina17 the first step you need get what will be the common denominator - ok. ?

jhonyy9 (jhonyy9):

comen please i like learn you how you can solve same exercises

jhonyy9 (jhonyy9):

do you like it ?

OpenStudy (lina17):

i want to know how to do it step by step

jhonyy9 (jhonyy9):

so than answer this my first question please

OpenStudy (lina17):

how do i get a common denominator

OpenStudy (anonymous):

well sometimes the fastest way to get a common denominator is to multiply the denominators of the fractions in this case 5*3=15 so 15 is a common denominator

OpenStudy (lina17):

oh okay

jhonyy9 (jhonyy9):

4/5 x + 4/3 = 2x Bill but dont forget it the right side there is again one denominator - yes ?

OpenStudy (anonymous):

yeah sorry

OpenStudy (lina17):

so was that wrong

OpenStudy (anonymous):

no just there is a 2/1 multiplying the x and so for future reference jhonyy9 is pointing out that there could be a fraction on both sides of the equation. If this happens then you would multiply all of the denominators together

OpenStudy (lina17):

oh okay

OpenStudy (anonymous):

after multiplying by a common denominator all the fractions should be gone. In this case with 4/5x+4/3=2x what do you get when you multiply by the common denominator?